# Structure-preserving space discretization of differential and nonlocal constitutive relations for port-Hamiltonian systems

### v1

@@ Abstract
To illustrate our results, the $1$D nanorod case and the shear beam model are considered, which are given by differential and implicit constitutive relations for which a Stokes-Lagrange structure along with boundary energy ports naturally occur.
Then, these results are extended to the nonlinear $2$D incompressible Navier-Stokes equations written in a vorticity–stream function formulation. It is first recast as a pH system defined with a Stokes-Lagrange structure along with a modulated Stokes-Dirac structure. A careful structure-preserving space discretization is then performed, leading to a finite-dimensional pH system. Theoretical and numerical results show that both enstrophy and kinetic energy evolutions are preserved both at the semi-discrete and fully-discrete levels.
@@ 1 Introduction
For infinite-dimensional pH systems, Stokes-Dirac structures have been introduced in [76], extending the in-domain port variables pairs with pairs of boundary port variables, which are necessary to write the usual (mass, energy, momentum, etc.) balance equations. Skew symmetry in the Hamiltonian differential operator, integration by parts, and Stokes theorem may be combined to obtain a canonical representation for these infinite-dimensional pH systems and define appropriate pairs of boundary port variables in which physically consistent boundary conditions, boundary control and observation variables may be expressed. In the linear case, the semigroup approach applied to distributed parameter linear pH systems allows a parametrization of all admissible boundary conditions (using these pairs of boundary port variables), which guarantee well-posedness of the resulting partial differential equations (PDE) (see [53] for the 1D case or [14] for the $n$D case). The literature on distributed pH systems has recently grown considerably. Both theoretical developments and application-based studies are surveyed in [65]. Recently, pH models for distributed parameter systems (either linear or not) have been proposed and investigated in various application domains, ranging from flexible mechanics (see [64] for a systematic methodology for pH modeling of flexible systems) to fluid dynamics (see for instance [18] for a literature review on pH modeling in fluid dynamics), electromagnetism [31], thermo-magneto-hydronamics (see for instance [78], for the pH model of tokamak plasma dynamics), and phase separation [77, 7], just to cite a few examples.
@@
In most of these applications (and previously cited papers), the boundary pH formulation relies on the assumption that the Hamiltonian differential operator, together with a generating Hamiltonian functional, are explicitly known and only depend on the state variables and not on their spatial derivatives. However, constraints may stem from implicit energy constitutive equations, affecting the relationship between effort and energy state variables. In this case, pH systems may be defined on a Lagrangian subspace (see [74]) or submanifold (for the nonlinear case, see [75]), as recently introduced in the finite-dimensional case. The energy is then no longer defined by a function on the state space, but instead by reciprocal constitutive relations between the state and co-state variables. Stokes-Lagrange subspaces have been proposed to cope with the infinite-dimensional case (see [50] or [55] for 1D formulation and [6] for $n$D extension of this idea). It is then possible to represent systems where the Hamiltonian functional is constrained, implicitly defined and/or depends on spatial derivatives of the state. Examples of pH systems defined on Stokes-Lagrange subspaces include the implicit formulation of the Allen-Cahn equation [81], the Dzektser equation governing the seepage of underground water (see [28] for the original model or [6, 45] for pH formulations), or nonlocal visco-elastic models for nanorods [42, 43]. The authors have also shown how constrained pH systems defined on Stokes-Lagrange subspaces may formalize the reduction of energy constitutive equations, considering the simplification from the Reissner-Mindlin thick plate model to the Kirchhoff-Love thin plate model, and the low-frequency approximation of Maxwell’s equations, as two examples of such a reduction [6].
@@
Regarding the structure-preserving spatial discretization of pH systems, various approaches have been proposed, including finite differences on staggered grids [72], pseudo-spectral methods [60, 79], methods based on the finite discrete exterior calculus [71, 44] and finite element methods. Among these approaches, only finite element methods have been successfully applied to a large number of different 2D and 3D examples, either linear or not, with non-trivial geometries (in that particular sense, they may be regarded as more general). Finite elements have been introduced for the structure-preserving discretization of 1D transmission line problems in [36]. The idea was then generalized for non linear shallow water equations in [63]. These methods are mixed finite element methods (MFEM) such as those analyzed in [10], but extended for pH systems with time-varying boundary energy flows. Different directions have been followed to generalize these early 1D MFEM results. In [49], structure-preserving is obtained by choosing different compatible finite element spaces for the effort and flow variables approximations in the weak formulation of pH systems. In [17] (generalized from [82]), the authors define a virtual and redundant dual field representation of the original pH system to obtain the finite dimensional without the need to discretize the Hodge metric in the constitutive equations. Finally, in [20], a partitioned finite element method (PFEM) is proposed, based on the integration by parts of some of the balance equations, chosen to make the boundary control appear and obtain a finite-dimensional square invertible pH matrix representation with sparse matrices. We propose in this paper an extension of the PFEM to pH systems defined on Stokes-Lagrange subspaces.
@@
The PFEM is defined and applied to shallow water equations (in 1D and 2D), together with various extensions examples in [20]. Since then, it has been applied to many different application examples, either in 2D or 3D, and has been proven to preserve the central Dirac interconnection structure of pH systems (hence, the power continuity). Among these examples, one can cite [70] for the anisotropic heterogeneous $n$D heat equation, [8, 7] for the Cahn-Hillard and Allen-Cahn equations, [40] for the Maxwell equations, [16] for von Karman plate models, [13] for a linear thermo-elasticity model (through the interconnection of elasto-dynamics and heat equations), [12] for Mindlin and Kirchhoff plate models, [39] for the incompressible Navier-Stokes equation, [19] for 2D shallow water equations, or [18] for a survey on applications of the PFEM to fluid models. The optimal choice of finite elements for the PFEM has been analyzed for the 2D wave equation on different geometries and asymptotics for the error are given in [41]. An open simulation Python package for multiphysics pH systems relying on the PFEM is available at https://g-haine.github.io/scrimp/, with various examples. This package is further discussed with the help of some mechanical and thermodynamical examples in [32]. Note also that a benchmark of pH numerical models with a repository of structure-preserving discretization methods may be found at https://algopaul.github.io/PortHamiltonianBenchmarkSystems.jl/.
@@
This paper deals with the structure-preserving discretization of $n$D pH systems with implicitly defined energy (i.e., with differential constitutive equations for the energy). They are represented with two geometrical interconnection structures: a Stokes-Dirac structure for power continuity (balance equations) and a Stokes-Lagrange structure for the energy constitutive equations. Until now, to the best of our knowledge, the PFEM discretization of pH system with implicitly defined energy has been applied only in [5], for a simple 1D nanorod viscoelastic example. In this work, we propose to extend these early results to more fundamental 1D and 2D examples, where the nonlocal constitutive equations play a fundamental role, to discretize these examples and analyze numerical consequences, of both the implicit pH formulation and the structure preservation, on simulation numerical results.
@@ 2.1 Finite-dimensional case
Let us denote by $(\alpha,e)\in\mathbb{R}^{2n}$ the state and costate of our system, resp., $(f_{R},e_{R})\in\mathbb{R}^{2r}$ the resistive port, $(u_{D},y_{D})\in\mathbb{R}^{2n_{D}}$ the power control port, and $(u_{L},y_{L})\in\mathbb{R}^{2n_{L}}$ the energy control port. We dropped the time dependency to improve readability, as our primary focus is on linear time-invariant systems.
@@
$\begin{pmatrix}\frac{\textup{d}}{\textup{d}t}\alpha\\ f_{R}\\ -y_{D}\end{pmatrix}$ $=J\begin{pmatrix}e\\ e_{R}\\ u_{D}\end{pmatrix}$ with $\text{with }J=-J^{\top},$ (1a)
@@
$\begin{pmatrix}\frac{\textup{d}}{\textup{d}t}\alpha\\ f_{R}\\ -y_{D}\end{pmatrix}$ $=J\begin{pmatrix}e\\ e_{R}\\ u_{D}\end{pmatrix}$ with $\text{with }J=-J^{\top},$ (2a)
$e$ $=Q\alpha$ with $\text{with }Q=Q^{\top}>0,$ (2b)
$e_{R}$ $=R\,f_{R}$ with $\text{with }R=R^{\top}\geq 0.$ (2c)
@@
Let $P$ and $S$ as in Definition 1. Let $\alpha,e\in\mathbb{R}^{n}$ and $y_{L},u_{L}\in\mathbb{R}^{n}_{L}$, satisfying (1b). Then, there exists a unique pair $(\lambda,\tilde{u}_{L})\in\mathbb{R}^{n+n_{L}}$ such that:
@@
Let $J$, $P$, $S$ and $R$ as in Definition 1. Decomposing $P=\begin{bmatrix}P_{1,1}&P_{1,2}\\ P_{2,1}&P_{2,2}\end{bmatrix}$ into blocks of appropriate sizes, the associated linear pH system (1) written in latent variables reads:
@@
Remark 4.
In this setting, the symmetry of $P^{*}S$ is expressed with the usual Euclidean inner product on $\mathbb{R}^{n+n_{L}}$, hence $P^{*}=P^{\top}$, and checking the symmetry of the operator $P^{*}S$ amounts to checking the symmetry of the matrix $P^{\top}S$. However, when applying the Finite Element Method, an invertible mass matrix $M$ appears on the left-hand side of (4b). The symmetry condition is then expressed using the weighted inner product $\langle\cdot,\cdot\rangle_{M^{-1}}$, and checking the symmetry of $P^{*}S$ is then achieved by checking the symmetry of the matrix $P^{\top}M^{-1}S$.

### v2

@@ Abstract
To illustrate our results, the 11D nanorod case and the shear beam model are considered, which are given by differential and implicit constitutive relations for which a Stokes-Lagrange structure along with boundary energy ports naturally occur.
Then, these results are extended to the nonlinear 22D incompressible Navier-Stokes equations written in a vorticity–stream function formulation. It is first recast as a pH system defined with a Stokes-Lagrange structure along with a modulated Stokes-Dirac structure. A careful structure-preserving space discretization is then performed, leading to a finite-dimensional pH system. Theoretical and numerical results show that both enstrophy and kinetic energy evolutions are preserved both at the semi-discrete and fully-discrete levels.
@@ 1 Introduction
For infinite-dimensional pH systems, Stokes-Dirac structures have been introduced in [76], extending the in-domain port variables pairs with pairs of boundary port variables, which are necessary to write the usual (mass, energy, momentum, etc.) balance equations. Skew symmetry in the Hamiltonian differential operator, integration by parts, and Stokes theorem may be combined to obtain a canonical representation for these infinite-dimensional pH systems and define appropriate pairs of boundary port variables in which physically consistent boundary conditions, boundary control and observation variables may be expressed. In the linear case, the semigroup approach applied to distributed parameter linear pH systems allows a parametrization of all admissible boundary conditions (using these pairs of boundary port variables), which guarantee well-posedness of the resulting partial differential equations (PDE) (see [53] for the 1D case or [14] for the nnD case). The literature on distributed pH systems has recently grown considerably. Both theoretical developments and application-based studies are surveyed in [65]. Recently, pH models for distributed parameter systems (either linear or not) have been proposed and investigated in various application domains, ranging from flexible mechanics (see [64] for a systematic methodology for pH modeling of flexible systems) to fluid dynamics (see for instance [18] for a literature review on pH modeling in fluid dynamics), electromagnetism [31], thermo-magneto-hydro dynamics (see for instance [78], for the pH model of tokamak plasma dynamics), and phase separation [77, 7], just to cite a few examples.
@@
In most of these applications (and previously cited papers), the boundary pH formulation relies on the assumption that the Hamiltonian differential operator, together with a generating Hamiltonian functional, are explicitly known and only depend on the state variables and not on their spatial derivatives. However, constraints may stem from implicit energy constitutive equations, affecting the relationship between effort and energy state variables. In this case, pH systems may be defined on a Lagrangian subspace (see [74]) or submanifold (for the nonlinear case, see [75]), as recently introduced in the finite-dimensional case. The energy is then no longer defined by a function on the state space, but instead by reciprocal constitutive relations between the state and co-state variables. Stokes-Lagrange subspaces have been proposed to cope with the infinite-dimensional case (see [50] or [55] for 1D formulation and [6] for nnD extension of this idea). It is then possible to represent systems where the Hamiltonian functional is constrained, implicitly defined and/or depends on spatial derivatives of the state. Examples of pH systems defined on Stokes-Lagrange subspaces include the implicit formulation of the Allen-Cahn equation [81], the Dzektser equation governing the seepage of underground water (see [28] for the original model or [6, 45] for pH formulations), or nonlocal visco-elastic models for nanorods [42, 43]. The authors have also shown how constrained pH systems defined on Stokes-Lagrange subspaces may formalize the reduction of energy constitutive equations, considering the simplification from the Reissner-Mindlin thick plate model to the Kirchhoff-Love thin plate model, and the low-frequency approximation of Maxwell’s equations, as two examples of such a reduction [6].
@@
Regarding the structure-preserving spatial discretization of pH systems, various approaches have been proposed, including finite differences on staggered grids [72], pseudo-spectral methods [60, 79], methods based on the finite discrete exterior calculus [71, 44] and finite element methods. Among these approaches, only finite element methods have been successfully applied to a large number of different 2D and 3D examples, either linear or not, with non-trivial geometries (in that particular sense, they may be regarded as more general). Finite elements have been introduced for the structure-preserving discretization of 1D transmission line problems in [36]. The idea was then generalized for non linear shallow water equations in [63]. These methods are mixed finite element methods (MFEM) such as those analyzed in [10], but extended for pH systems with time-varying boundary energy flows. Different directions have been followed to generalize these early 1D MFEM results. In [49], structure-preserving is obtained by choosing different compatible finite element spaces for the effort and flow variables approximations in the weak formulation of pH systems. In [17] (generalized from [82]), the authors define a virtual and redundant dual field representation of the original pH system to obtain the finite dimensional without the need to discretize the Hodge metric in the constitutive equations. Finally, in [20], a partitioned finite element method (PFEM) is proposed, based on the integration by parts of some of the balance equations, chosen to make the boundary control appear and obtain a finite-dimensional square invertible pH matrix representation with sparse matrices. Note that the PFEM is a particular instance of the MFEM applied to pH systems, where the partition of the variables is a natural feature. We propose in this paper a new application of the PFEM to pH systems defined on Stokes-Lagrange subspaces.
@@
The PFEM is defined and applied to shallow water equations (in 1D and 2D), together with various extensions examples in [20]. Since then, it has been applied to many different application examples, either in 2D or 3D, and has been proven to preserve the central Dirac interconnection structure of pH systems (hence, the power continuity). Among these examples, one can cite [70] for the anisotropic heterogeneous nnD heat equation, [8, 7] for the Cahn-Hillard and Allen-Cahn equations, [40] for the Maxwell equations, [16] for von Karman plate models, [13] for a linear thermo-elasticity model (through the interconnection of elasto-dynamics and heat equations), [12] for Mindlin and Kirchhoff plate models, [39] for the incompressible Navier-Stokes equation, [19] for 2D shallow water equations, or [18] for a survey on applications of the PFEM to fluid models. The optimal choice of finite elements for the PFEM has been analyzed for the 2D wave equation on different geometries and asymptotics for the error are given in [41]. An open simulation Python package for multiphysics pH systems relying on the PFEM is available at https://g-haine.github.io/scrimp/, with various examples. This package is further discussed with the help of some mechanical and thermodynamical examples in [32]. Note also that a benchmark of pH numerical models with a repository of structure-preserving discretization methods may be found at https://algopaul.github.io/PortHamiltonianBenchmarkSystems.jl/.
@@
This paper deals with the structure-preserving discretization of nnD pH systems with implicitly defined energy (i.e., with differential constitutive equations for the energy). They are represented with two geometrical interconnection structures: a Stokes-Dirac structure for power continuity (balance equations) and a Stokes-Lagrange structure for the energy constitutive equations. Until now, to the best of our knowledge, the PFEM discretization of pH system with implicitly defined energy has been applied only in [5], for a simple 1D nanorod viscoelastic example. In this work, we propose to extend these early results to more fundamental 1D and 2D examples, where the nonlocal constitutive equations play a fundamental role, to discretize these examples and analyze numerical consequences, of both the implicit pH formulation and the structure preservation, on simulation numerical results.
@@ 2.1 Finite-dimensional case
Let us denote by $(\alpha,e)\in\mathbb{R}^{n}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\times\mathbb{R}^{n}}$ the state and costate of our system, resp., $(f_{R},e_{R})\in\mathbb{R}^{r}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\times\mathbb{R}^{r}}$ the resistive port, $(u_{D},y_{D})\in\mathbb{R}^{n_{D}}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\times\mathbb{R}^{n_{D}}}$ the power control port, and $(u_{L},y_{L})\in\mathbb{R}^{n_{L}}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\times\mathbb{R}^{n_{L}}}$ the energy control port [^1] Indices stand for Resistive, Dirac and Lagrange.. We dropped the time dependency to improve readability, as our primary focus is on linear time-invariant systems.
@@
$\begin{pmatrix}\frac{\textup{d}}{\textup{d}t}\alpha\\ f_{R}\\ -y_{D}\end{pmatrix}$ $=J\begin{pmatrix}e\\ e_{R}\\ u_{D}\end{pmatrix}$ $\text{with }J=-J^{\top},$ (1a)
@@
$\begin{pmatrix}\frac{\textup{d}}{\textup{d}t}\alpha\\ f_{R}\\ -y_{D}\end{pmatrix}$ $=J\begin{pmatrix}e\\ e_{R}\\ u_{D}\end{pmatrix}$ $\text{with }J=-J^{\top},$ (2a)
$e$ $=Q\alpha$ $\text{with }Q=Q^{\top}>0,$ (2b)
$e_{R}$ $=R\,f_{R}$ $\text{with }R=R^{\top}\geq 0.$ (2c)
@@
Let $P$ and $S$ as in Definition 1. Let $\alpha,e\in\mathbb{R}^{n}$ and $y_{L},u_{L}\in\mathbb{R}^{n_{L}}$, satisfying (1b). Then, there exists a unique pair $(\lambda,\tilde{u}_{L})\in\mathbb{R}^{n}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\times\mathbb{R}^{n_{L}}}$ such that:
@@
Let $J$, $P$, $S$ and $R$ as in Definition 1. Decomposing $P=\begin{bmatrix}P_{1,1}&P_{1,2}\\ P_{2,1}&P_{2,2}\end{bmatrix}$ into blocks of appropriate sizes (precisely, $P_{1,1}\in\mathbb{R}^{2n}$, $P_{2,2}\in\mathbb{R}^{2n_{L}}$), the associated linear pH system (1) written in latent variables reads:
@@
2.2 A toy model: a moving mass
In order to fix ideas and vocabulary, let us construct a simple worked out example.
The classical [73] port-Hamiltonian formalism in the linear case, i.e., with a quadratic Hamiltonian $H(\alpha):=\frac{1}{2}\alpha^{\top}Q\alpha$, reads:
$\left\{\begin{array}[]{ll}\frac{\textup{d}}{\textup{d}t}\alpha&=JQ\alpha+Bu_{D},\\ y_{D}&=B^{\top}Qe.\end{array}\right$.
where $\alpha$ is the energy variable (the variable of the Hamiltonian which often represents an energy), $J$ is the interconnection matrix (because it interconnects the elements of $e:=\textbf{{grad}}H=Q\alpha$), it is skew-adjoint and its graph222more precisely the graph of its inverse, see [73, Exercise 1, p. 17] is the Dirac structure the trajectories belong to, and $Q$ is a positive-definite self-adjoint matrix that relates (through constitutive relations) energy variables to co-energy variables $e=Q\alpha$ (i.e., the derivative of the Hamiltonian with respect to $\alpha$). Finally, $B$ is a control matrix, and $u_{D}$ and $y_{D}$ are the control (or input) and, respectively, the observation (or output). Together, $u_{D}$ and $y_{D}$ constitutes the power control port as it allows to control the behavior of the power (because the Hamiltonian is thought as an energy) flowing into the system. Indeed, the power balance is trivially:
$\frac{\textup{d}}{\textup{d}t}H=y_{D}^{\top}u_{D}.$
Now, consider a mass $m$ moving along a fixed $x$-axis, which can be controlled according to Newton’s second law: $F=ma$, with $F$ the applied force and $a$ the resulting acceleration. We argue that such a simple system does not fit into the classical framework briefly recalled above. Indeed, the best that can be done is to get $Q$ being only a non-negative self-adjoint matrix. Lagrangian subspace allows us to go further to include this simple example.
Let us denote $q$ the position of the mass on the $x$-axis, $p:=m\dot{q}$ its linear momentum, with $\dot{q}$ its velocity. The kinetic energy gives a Hamiltonian $H:=\frac{p^{2}}{2m}$. Taking the position and the linear momentum as energy variables (as usual in Hamiltonian mechanics), the co-energy variables are respectively $e_{q}:=0$ and $e_{p}:=q$. Clearly, the problem lies in the $q$-independence of the Hamiltonian. Nevertheless, one can write:
$\frac{\textup{d}}{\textup{d}t}\underbrace{\begin{pmatrix}q\\ p\end{pmatrix}}_{\alpha}=\underbrace{\begin{bmatrix}0&1\\ -1&0\end{bmatrix}}_{J}\underbrace{\begin{pmatrix}0\\ \dot{q}\end{pmatrix}}_{e}+\underbrace{\begin{bmatrix}0\\ 1\end{bmatrix}}_{B}\underbrace{F}_{u}.$
Now, observe that:
$\begin{pmatrix}\alpha\\ e\end{pmatrix}=\begin{pmatrix}q\\ m\dot{q}\\ 0\\ \dot{q}\end{pmatrix}=\begin{bmatrix}1&0\\ 0&1\\ 0&0\\ 0&\frac{1}{m}\end{bmatrix}\begin{pmatrix}q\\ m\dot{q}\end{pmatrix}=\begin{bmatrix}P\\ S\end{bmatrix}\underbrace{\begin{pmatrix}z_{q}\\ z_{p}\end{pmatrix}}_{z},$
and the $z$ variables are now the latent variables that allow to recover the Hamiltonian $H=\frac{1}{2}z^{\top}P^{\top}Sz$ from the matrices $P$ and $S$, which define a Lagrangian subspace thanks to the following properties: $P^{\top}S=S^{\top}P$ and $\text{\rm rank}\begin{bmatrix}P\\ S\end{bmatrix}=2$.
Altogether, this guarantees that the moving mass is a port-Hamiltonian system in the sense of Definition 1 (without resistive port, nor energy control port).
Assume now a coefficient $k\geq 0$ modeling a friction proportional to the velocity of the mass, i.e., a friction force $F_{R}:=-k\dot{q}$. Then, such a dissipative term can be added to the previous model by adding a resistive port $(f_{R},e_{R})$ as follows:
$\begin{pmatrix}\frac{\textup{d}}{\textup{d}t}\alpha\\ f_{R}\end{pmatrix}=\begin{bmatrix}J&\begin{matrix}0\\ -1\end{matrix}\\ \begin{matrix}0&1\end{matrix}&0\end{bmatrix}\begin{pmatrix}e\\ e_{R}\end{pmatrix}+\begin{bmatrix}0\\ 1\\ 0\end{bmatrix}F,\quad e_{R}=kf_{R}.$
Finally, assume now that we want to observe the position of the mass rather than its force or velocity (e.g., because this output is needed for interconnection with another system); it can be achieved by the introduction of an energy control port (as it controls the behavior of the Hamiltonian directly instead of its time-derivative), as follows. Let us define $(y_{L},u_{L})$ the energy control port by:
$\begin{pmatrix}\alpha\\ u_{L}\\ e\\ y_{L}\end{pmatrix}=\begin{pmatrix}q\\ m\dot{q}\\ u_{L}\\ 0\\ \dot{q}\\ y_{L}\end{pmatrix}=\begin{bmatrix}1&0&0\\ 0&1&0\\ 0&0&1\\ 0&0&1\\ 0&\frac{1}{m}&0\\ 1&0&0\end{bmatrix}\begin{pmatrix}q\\ m\dot{q}\\ \tilde{u}_{L}\end{pmatrix}=\begin{bmatrix}\tilde{P}\\ \tilde{S}\end{bmatrix}\underbrace{\begin{pmatrix}z_{q}\\ z_{p}\\ \tilde{u}_{L}\end{pmatrix}}_{\tilde{z}}$,
where $\tilde{u}_{L}$ is the latent control. Then, a Hamiltonian reads:
$\tilde{H}=\frac{1}{2}\tilde{z}^{\top}\tilde{P}^{\top}\tilde{S}\tilde{z}=\frac{1}{2}z^{\top}P^{\top}Sz+\frac{1}{2}z_{q}^{\top}\tilde{u}_{L},$
and the associated power balance becomes:
$\frac{\textup{d}}{\textup{d}t}\tilde{H}=-kf_{R}^{2}+y_{D}^{\top}u_{D}+\dot{y}_{L}^{\top}u_{L}.$
Beside its simplicity, this example motivates the energetic and modular approach provided by Dirac structure and Lagrangian subspace. Furthermore, this geometric standpoint allows one to switch input and output. For instance, one could control the position of the mass in the previous example by only switching the roles of $u_{L}$ and $y_{L}$. The price to pay is that the control becomes a constraint that has to be satisfied, e.g., using a Lagrange multiplier, for the resolution.
In summary, Dirac structures are meant for power-preserving exchanges between each elements of the system (through the interconnection matrix), while Lagrange structure enforce Maxwell’s reciprocity conditions for the constitutive relations ($P^{\top}e=S^{\top}\alpha$), from which a Hamiltonian satisfying the expected power balance can be derived, thanks to the latent variables.